OpenStudy (ria23):

Urgent The temperature on one day in January for Phoenix, AZ, is 52 degrees Fahrenheit. On the same day in Barrow, AK, the temperature is -23 degrees Fahrenheit. what is the difference between the two temperatures?

OpenStudy (ria23):

I thought it was just 52-23, now I'm second guessing, is it 52-(-23)?

OpenStudy (ria23):

@whpalmer4

OpenStudy (ria23):

@johnweldon1993 @shiraz14 @tgawade

OpenStudy (anonymous):

Yes second one is correct !! 52 - (-23) = -----------?

OpenStudy (whpalmer4):

Don't forget that when you subtract a negative number, it is the same as adding the positive number with the same magnitude: \[5-(-3) = 5 + 3 = 8\] If you think of subtraction of a positive number as moving that number of spaces to the left on a number line, subtraction of a negative number is simply moving that number of spaces in the other direction, or to the right.

OpenStudy (ria23):

Thank yhu.